# Circumventing shortcuts in audio-visual deepfake detection datasets with unsupervised learning

### v1

@@ Abstract
Good datasets are essential for developing and benchmarking any machine learning system. Their importance is even more extreme for safety critical applications such as deepfake detection—the focus of this paper. Here we reveal that two of the most widely used audio-video deepfake datasets suffer from a previously unidentified spurious feature: the leading silence. Fake videos start with a very brief moment of silence and based on this feature alone, we can separate the real and fake samples almost perfectly. As such, previous audio-only and audio-video models exploit the presence of silence in the fake videos and consequently perform worse when the leading silence is removed. To circumvent latching on such unwanted artifact and possibly other unrevealed ones we propose a shift from supervised to unsupervised learning by training models exclusively on real data. We show that by aligning self-supervised audio-video representations we remove the risk of relying on dataset-specific biases and improve robustness in deepfake detection.
@@ 1 Introduction
The progress in automated deepfake detection is fueled by the datasets developed by the research community. Good quality data is essential for both training and benchmarking the progress of these methods. Over the last years, numerous datasets have been proposed. Among the audio-video deepfake datasets, there have been released datasets that alter both streams [22, 6] or only one [54, 33, 43] ; datasets with full-video [22] or only local [6, 4] manipulations.
However, care has to be taken to ensure a good deepfake detection dataset. Any asymmetry in the preparation of fakes and reals can result in biases that correlate spuriously with the groundtruth label. For example, in the image domain, different preprocessing pipelines [8] or types of resizing [40] between real and fake samples paint an overly optimistic picture. In the audio domain, the very popular ASVSpoof19 dataset [54] has been shown to leak information about the labels in the form of silence duration [32] or bitrate information [3]. Since many of the deepfake detection models are high capacity, they can easily use such artifacts as shortcuts for learning, consequently greatly impacting their ability to generalize in a real case scenario.
@@
To summarize, our work makes the following contributions: 1. We expose a previously unknown spurious feature in two of the most widely adopted deepfake detection datasets. 2. We analyze the impact of this shortcut on various state-of-the-art models. 3. We show that a way to mitigate such shortcuts is by training on real data only and we introduce a new method in this direction.
@@ 2 Related work
Audio-video deepfake detection. Many approaches for deepfake detection on videos have focused on the visual stream of information only [1, 17, 60, 18, 20, 47, 5]. But with the recent introduction of audio-visual datasets, (FakeAVCeleb [22], Deepfake TIMIT [24, 44], KODF [26], LAV-DF [4], AV-Deepfake-1M [6]), more research has shifted towards models that exploit both audio and video cues [16, 36, 45, 19, 57]. An emerging trend in this direction is the use of pretrained representations in a self-supervised way [16, 36]. But, different from our approach, these methods train the representations from scratch and use them as a first step in a more elaborated pipeline: supervised classification [36] or anomaly detection [16]. There are also works that similarly to us exploit the pretrained audio-video AV-HuBERT model [46] to extract representations [45, 19], but all of those methods are trained in the fully supervised learning paradigm.
Unsupervised deepfake detection. To make detection more generalizable across generators, a new direction is to depart from the supervised paradigm and resort only on real data. There are two main classes of such unsupervised approaches: methods that rely on consistency checks and methods that treat the problem as an anomaly detection one. Among the first class, prior work proposed to verify that the audio and visual streams align at a semantic level, for example from the point of view of spoken content [28, 2], or at a representation level, based on the alignment of audio and video features [16, 41]. Consistency checks have also been used with respect to the identity of a speaker (comparing a query sample against real audio [38] or real images [41] of the target speaker) or between image and text modalities [41]. For the second class, deepfake detection as anomaly detection, recent work in the image domain has used the reconstruction loss of the query image to tell whether it is anomalous [42, 14]; for example, Ricker et al. [42] make the observation that images generated by a latent diffusion model (LDM) are easier to reconstruct by the LDM than real images. For the audio-visual domain, Feng et al. [16] use both classes of approaches: they use consistency checks to estimate synchronization between the two streams and then flag anomalies using density estimation.
@@ 3 Silence bias in audio-video datasets
In this section, we show that two popular datasets (Sec. 3.1) have a silence bias. We analyze its behavior and show that a simple classifier based on the leading silence can obtain almost perfect separation between fake and real samples (Sec. 3.2). This implies that the performance of prior work is susceptible to have been overestimated. For this reason we analyze its impact on various audio and audio-visual methods (Sec. 3.3).
@@ 3.1 Datasets
FakeAVCeleb contains 500 real videos from VoxCeleb2 and 19.5k fake videos (10k FVFA, 9k FVRA, 500 RVFA). The fake visual content was generated with face swapping methods (Faceswap [25] and FSGAN [34]) or the Wav2Lip lip syncing approach [39]. The fake audio content was generate with the voice cloning tool SV2TTS [21]. The dataset is diverse across age groups, genders, races, as well as with respect to the number of subjects in a single video, their placement, the visual and audio quality.
AV-Deepfake1M is a large scale dataset, which consists of over one million videos and 2k subjects. As opposed to the FakeAVCeleb dataset, here the manipulations are local and consists of word-level replacements, insertions and deletions. The text manipulations are generated with the ChatGPT large language model. The fake video content is generated with the lip syncing method TalkLip [53], while the fake audio content is generated with the VITS [23] or YourTTS [7] methods. The authors ensure that the synthesized words share the same background noise with the full audio, by first extracting the audio noise with the Denoiser method [15] and then adding it to the synthesized words. The dataset is split into train, validation and test, with the test having a different set of speakers than those encountered at the train and validation splits.
@@ 3.2 Analysis of leading silence
We start by analyzing the silence distribution of the real and fake samples in the two considered datasets. We define the duration of the leading silence as the moment when the magnitude of the audio exceeds a certain threshold $\tau$. For this experiment, we select this threshold to be $5\cdot 10^{-4}$, but as we will shortly see, the results are robust to its choice. We carry this analysis only on the real (RVRA) and fully fake (FVFA) videos. For the AV-Deepfake1M dataset we use the the validation split, while for the FakeAVCeleb dataset we pick 30% random videos (since this dataset does not have a standard split).
The results are shown in Figure 2. We observe that the real videos start with noise, while the fake samples have a leading silence of around 25–30 ms. The duration of silence of fake sample is similar for both datasets, although the distribution is much sharper for the AV-Deepfake1M dataset. If we were to rank the samples based on this feature we would obtain an area under curve of the receiver operating characteristic curve (AUC) of over 98% for either datasets.
@@
Why do fake samples have a leading silence? Given that we do not have access to generation process of the two datasets, it is challenging to pinpoint the exact reason for the occurrence of the leading silence bias. However, we speculate that this happens when the audio may be slightly shorter than the video counterpart. Note that this is a different reason from the silence observed in audio only datasets [32]. There the real audios had a leading silence, while the synthesized speech was silence-free. In the case of audio-video datasets it might be challenging to completely avoid this problem, but an easy solution (for this particular bias) is to trim the start of both the real and fake samples. This is what we do in the next experiments.
@@ 3.3 Impact on prior work
  • MDS [10] (modality dissonance score), which is an audio-visual method that estimates the mismatch between audio and video segments along a video. The score is computed as the distance between audio and visual features. The audio features are extracted with a CNN from Mel-frequency cepstral coefficients (MFCCs), while the visual features use a 3D ResNet.
  • AVAD [16] (audio-visual anomaly detection), which is an audio-visual method trained on real data only. The approach has two steps: first, it estimates the desynchronization between audio and video; then it estimates whether these patterns are typical of real data or anomalous (indicating a fake). The method employs a 2D$+$3D ResNet-18 on the vision side and a VGG-M on Mel spectrograms on the audio side.
We train the RawNet2 and MDS methods on both datasets using the code provided by the authors. For the AVAD we use the provided checkpoint (trained on LRS [49]) and do not retrain it since training code is not available. We evaluate the methods in two settings: on the corresponding evaluation set and on a trimmed version of the same evaluation set. The trimming is done by removing the first 40 ms from all samples. Since the datasets use 25 frames per second, the trimming process removes the first frame. For training we do not perform trimming, but use the original dataset. As a baseline we include our silence classifier described in the previous section.
The results are shown in Table 1. First, we observe that the audio-based RawNet2 is the top performing method. Moreover, we see that the silence classifier is better than two other approaches: MDS and AVAD. The relatively worse performance of AVAD indicates that this method dose not latch on the silence information.
In terms of the impact of the leading silence, we see that when removing this information by trimming, the methods are affected differently. As expected, the silence classifier is affected the most and its performance drops down to random chance on the trimmed data. RawNet2 is not as affected on the FakeAVCeleb, presumably because there is still enough information throughout the signal, but it suffers a larger hit on the more challenging AV-DeepFake1M, which has only partially-manipulated samples. MDS is affected by the leading silence bias on both FakeAVCeleb and AV-Deepfake1M, though significantly stronger on AV-Deepfake1M. AVAD is the most robust method, since this approach does not explicitly model the silence information. Finally, we notice that the RawNet2 audio-only method is best even on the trimmed settings. This is noteworthy since it is the first time an audio-only method has been applied to these datasets. This suggests that the audio information is an important source of information that is often overlooked by prior work.
@@
Method Mod Trim: Trim: Trim: Trim:
@@ 4 Modeling real data for deepfake detection
The previous section indicated that audio-visual datasets exhibit a silence bias which can easily be exploited. We want to develop a method that is robust to this bias (and possibly other uncovered ones) and still performs well. We have seen that models trained only on real data [16] are promising in being robust to the silence shortcut, but the performance was modest. To further improve them we propose to build on top of audio-focused self-supervised features. Self-supervised features have shown strong generalization for both visual [35, 13] and audio [37, 38] deepfake detection. We choose audio-focused self-supervised features because they showed strong performance (as indicated the previous section). Note that we cannot rely on audio-only models because there are cases where manipulations appear only in the visual domain (the fake video, real audio case).
Figure 4: Overview of the AVH-Align method. A: We use the pretrained AV-HuBERT model to extract self-supervised features which we further align with a learnable network $\Phi$. B: At training we maximize the alignment score $\Phi_{ii}$, between the audio features $\mathbf{a}_{i}$ at time step $i$ and the corresponding video features $\mathbf{v}_{i}$, while minimizing the alignment $\Phi_{ik}$ to the other features $\mathbf{v}_{k}$ in a neighboring window $\mathcal{N}(i)$.
@@ 4.1 Method
Self-supervised features. We use AV-HuBERT to represent both the audio and visual content of a video. AV-HuBERT is a Transformer network trained in a self-supervised way to predict iteratively refined centroids from masked inputs. The features extracted by AV-HuBERT encode audio information as proved by its strong performance on tasks such lip reading or noisy audio-visual speech recognition. We extract the audio and visual representations independently: to extract audio features we mask the visual input, to extract visual features we mask the audio input. For a video we obtain audio features $\mathbf{a}_{i}$ and visual features $\mathbf{v}_{i}$ features for each time step $i$. Both representations are 1024 dimensional and have a temporal resolution of 25 frames per second.
@@
The MLP has four layers, which progressively reduce the feature dimensionality, with layers mapping from the AV-HuBERT feature size of 1024 to 512, 256, 128, and finally to a single output. Each hidden layer includes Layer Normalization and ReLU activation.
@@
where $\mathcal{N}(i)$ represents the temporal neighborhood around the frame $i$; in our case we choose the 30 neighboring frames around frame $i$. We define the loss across the entire video as the average:
@@
The final loss is similar to the contrastive loss InfoNCE [51] which was also used for deepfake detection [16, 36].
Inference. Once $\Phi$ is learned, we can estimate the alignment score $\Phi_{ii}$ for each audio-video frame pair in a video. Then we compute an overall alignment score for the entire video by pooling the per-frame scores using log-sum-exp. We estimate the fakeness score as the complement of the alignment probability (Eq. 2).
Supervised variant. To understand the impact of the silence bias on the standard supervised learning paradigm, we design a supervised variant, AVH-Align/sup, that uses the same features and alignment network as AVH-Align, but a classification loss. In this setup we assume that apart from the real videos, we also have access to fake videos in the training set, with corresponding labels $y$. We obtain a per-video alignment score by pooling the per-frame scores $\Phi_{ii}$ using the log-sum-exp function and optimize the binary cross-entropy (CE) loss:
$\mathcal{L}_{\text{sup}}=\text{CE}(\log\text{sum}\exp(\Phi_{ii}),y)$ (4)
FakeAVCeleb AV-Deepfake1M
Method Modality Train type Train data Trim: Trim: Trim: Trim:
@@
[…]
Table 2: Comparison of AVH-Align and AVH-Align/sup on FakeAVCeleb and AV-Deepfake1M original and trimmed datasets. AVH-Align is not impacted by the presence of leading silence in the original datasets, showing similar performance on the trimmed versions. AVH-Align/sup has a maximum degrading relative performance of $28.4\%$ when tested on AV-Deepfake1M trimmed compared to AV-Deepfake1M original.
@@
Refer to caption Refer to caption Refer to caption
Figure 5: Per frame fakeness probabilities for AVH-Align and AVH-Align/sup on AV-Deepfake1M. AVH-Align/sup always marks the first frame—corresponding to the leading silence—as fake, thus confirming that it uses the bias to distinguish between real and fake videos. AVH-Align is not affected by the presence of the leading silence. The fakeness probabilities for AVH-Align can be interpreted as misalignment probabilities, which is why they are higher during or after the manipulated region: the separation and re-composition of the audio and video streams causes misalignment throughout the tampered video.

### v3

@@ Abstract
Good datasets are essential for developing and benchmarking any machine learning system. Their importance is even more extreme for safety critical applications such as deepfake detection—the focus of this paper. Here we reveal that two of the most widely used audio-video deepfake datasets suffer from a previously unidentified spurious feature: the leading silence. Fake videos start with a very brief moment of silence and, on the basis of this feature alone, we can separate the real and fake samples almost perfectly. As such, previous audio-only and audio-video models exploit the presence of silence in the fake videos and consequently perform worse when the leading silence is removed. To circumvent latching on such an unwanted artifact and possibly other unrevealed ones, we propose a shift from supervised to unsupervised learning by training models exclusively on real data. We show that by aligning self-supervised audio-video representations we remove the risk of relying on dataset-specific biases and improve robustness in deepfake detection.
@@ 1 Introduction
The progress in automated deepfake detection is fueled by the datasets developed by the research community. Good quality data is essential for both training and benchmarking the progress of these methods. In recent years, numerous datasets have been proposed. Among the audio-video deepfake datasets, there have been released datasets that alter both streams [23, 6] or only one [56, 36, 46] ; datasets with full-video [23] or only local [6, 4] manipulations.
However, care has to be taken to ensure a good deepfake detection dataset. Any asymmetry in the preparation of fakes and reals can result in biases that correlate spuriously with the groundtruth label. For example, in the image domain, different preprocessing pipelines [8] or types of resizing [43] between real and fake samples paint an overly optimistic picture. In the audio domain, the very popular ASVSpoof19 dataset [56] has been shown to leak information about labels in the form of silence duration [35] or bitrate information [3]. Since many of the deepfake detection models are high capacity, they can easily use such artifacts as shortcuts for learning, consequently greatly impacting their ability to generalize in a real case scenario.
@@
To summarize, our work makes the following contributions: 1. We expose a previously unknown spurious feature in two of the most widely adopted deepfake detection datasets. 2. We analyze the impact of this shortcut on various state-of-the-art models. 3. We show that a way to mitigate such shortcuts is by training on real data only and we introduce a new method in this direction. Our code is available at: https://github.com/bit-ml/AVH-Align.
@@ 2 Related work
Audio-video deepfake detection. Many approaches for deepfake detection on videos have focused on the visual stream of information only [1, 18, 60, 19, 21, 50, 5]. But with the recent introduction of audio-visual datasets, (FakeAVCeleb [23], Deepfake TIMIT [25, 47], KODF [28], LAV-DF [4], AV-Deepfake-1M [6]), more research has shifted towards models that exploit both audio and video cues [17, 39, 48, 20, 58, 27]. An emerging trend in this direction is the use of pretrained representations in a self-supervised way [17, 39]. But, different from our approach, these methods train the representations from scratch and use them as a first step in a more elaborated pipeline: supervised classification [39] or anomaly detection [17]. There are also works that similarly to us exploit the pretrained audio-video AV-HuBERT model [49] to extract representations [48, 20], but all of those methods are trained in the fully supervised learning paradigm.
Unsupervised deepfake detection. To make detection more generalizable across generators, a new direction is to depart from the supervised paradigm and resort only to real data. There are two main classes of such unsupervised approaches: methods that rely on consistency checks and methods that treat the problem as an anomaly detection one. Among the first class, prior work proposed to verify that audio and visual streams align at a semantic level, for example, from the point of view of spoken content [29, 2], or at a representation level, based on the alignment of audio and video features [17, 44]. Consistency checks have also been used with respect to the identity of a speaker (comparing a query sample against real audio [41] or real images [44] of the target speaker) or between image and text modalities [44]. For the second class, deepfake detection as anomaly detection, recent work in the image domain has used the reconstruction loss of the query image to tell whether it is anomalous [45, 14]; for example, Ricker et al. [45] make the observation that images generated by a latent diffusion model (LDM) are easier to reconstruct by the LDM than real images. For the audio-visual domain, Feng et al. [17] use both classes of approaches: they use consistency checks to estimate synchronization between the two streams and then flag anomalies using density estimation.
@@ 3 Silence bias in audio-video datasets
In this section, we show that two popular datasets (Sec. 3.1) have a silence bias. We analyze its behavior and show that a simple classifier based on the leading silence alone can obtain almost perfect separation between fake and real samples (Sec. 3.2). This implies that the performance of prior work is susceptible to have been overestimated. For this reason we analyze its impact on various audio and audio-visual methods (Sec. 3.3).
@@ 3.1 Datasets
FakeAVCeleb contains 500 real videos from VoxCeleb2 and 19.5k fake videos (10k FVFA, 9k FVRA, 500 RVFA). The fake visual content was generated with face swapping methods (Faceswap [26] and FSGAN [37]) or the Wav2Lip lip syncing approach [42]. The fake audio content was generate with the voice cloning tool SV2TTS [22]. The dataset is diverse across age groups, genders, races, as well as with respect to the number of subjects in a single video, their placement, the visual and audio quality. In our experiments, we split the dataset in 70% (train and validation) and 30% (test). This split is kept consistent across all experiments.
AV-Deepfake1M is a large scale dataset, which consists of over one million videos and 2k subjects. As opposed to the FakeAVCeleb dataset, the manipulations here are local and consists of word-level replacements, insertions and deletions. The text manipulations are generated with the ChatGPT large language model. The fake video content is generated with the lip syncing method TalkLip [55], while the fake audio content is generated with the VITS [24] or YourTTS [7] methods. The authors ensure that the synthesized words share the same background noise with the full audio, by first extracting the audio noise with the Denoiser method [15] and then adding it to the synthesized words. The dataset is originally split into train, validation and test, with the test split having a different set of speakers than those encountered in the train and validation splits. For our experiments, we select training and validation samples from the original training split, and evaluate on 10k samples from the original validation split or on the full official test set.
@@ 3.2 Analysis of leading silence
We start by analyzing the silence distribution of the real and fake samples in the two considered datasets. We define the duration of the leading silence as the moment when the magnitude of the audio exceeds a certain threshold $\tau$. For this experiment, we select this threshold to be $5\cdot 10^{-4}$, but as we will shortly see, the results are robust to its choice. We carry this analysis only on the real (RVRA) and fully fake (FVFA) videos from each dataset’s test set.
The results are shown in Figure 2. We observe that the real videos start with noise, while the fake samples have a leading silence of around 25–30 ms. The silence duration of fake samples is similar for both datasets, although the distribution is much sharper for the AV-Deepfake1M dataset. If we were to rank the samples based on this feature we would obtain an area under curve of the receiver operating characteristic curve (AUC) of over 98% for either datasets.
@@
Why do fake samples have a leading silence? Given that we do not have access to generation process of the two datasets, it is challenging to pinpoint the exact reason for the occurrence of the leading silence bias. However, we speculate that this happens when the audio may be slightly shorter than the video counterpart. Note that this is a different reason from the silence observed in audio only datasets [35]. There the real audios had a leading silence, while the synthesized speech was silence-free. In the case of audio-video datasets it might be challenging to completely avoid this problem, but an easy solution (for this particular bias) is to trim the leading silence. This is what we do in the next experiments.
Biases in other datasets. We consider three more datasets: LAV-DF [4], AVLips [31], DFDC [16]. We find that LAV-DF has a slight trailing silence bias (65% AUC), but it does not exhibit the other two biases (leading silence and volume). In the case of AVLips and DFDC we did not identify any of the three biases. This may happen because the samples in these datasets have real audio and the unmodified audio matches the visual sequence more precisely. This is confirmed by a fine-grained analysis on FakeAVCeleb, where we observe that on the FVRA split the biases are diminished (62.4% AUC for lead silence), while they persist on the RVFA split (100% AUC for lead silence).
@@ 3.3 Impact on prior work
  • MDS [10] (modality dissonance score), which is an audio-visual method that estimates the mismatch between audio and video segments. The score is computed as the distance between audio and visual features.
  • AVAD [17] (audio-visual anomaly detection), which is an audio-visual method trained on real data only. The approach has two steps: first, it estimates the desynchronization between audio and video; then it estimates whether these patterns are typical of real data or anomalous.
We train the RawNet2 and MDS methods on both datasets using the code provided by the authors. For the AVAD we use the provided checkpoint (trained on LRS [52]) and do not retrain it since training code is not available. As a baseline we include our leading silence duration classifier described in the previous section.
We evaluate all methods in two settings: on the corresponding evaluation set and on a trimmed version of the same evaluation set. For the trimmed version, we discard the beginning of videos such that each of them starts with the first non-silent segment. We find the duration to discard by first computing the leading silence (using the same $5\cdot 10^{-4}$ threshold on the magnitude) and then rounding it up to the nearest multiple of 1 / FPS (the reciprocal of the video’s frame rate); this rounding ensures that the audio and video channels remain synchronized. For training, we do not perform trimming, but use the original dataset.
The results are shown in Table 1. First, we observe that the audio-based RawNet2 is the top performing method. Moreover, we see that the leading silence classifier is generally better than the two other approaches: MDS and AVAD. The relatively stable performance of AVAD indicates that this method dose not latch on the silence information.
In terms of the impact of the leading silence, we observe that removing it through trimming affects methods differently. As expected, the silence classifier is affected the most and its performance drops down to random chance on the trimmed data. RawNet2 is not as affected on the FakeAVCeleb, presumably because there is still enough information throughout the signal, but it suffers a larger hit on the more challenging AV-DeepFake1M, which has only partially-manipulated samples. MDS is affected by the leading silence bias on both FakeAVCeleb and AV-Deepfake1M, though significantly stronger on AV-Deepfake1M. AVAD is the most robust method, since this approach does not explicitly model the silence information. Finally, we notice that the RawNet2 audio-only method is best even on the trimmed settings. This is noteworthy since it is the first time an audio-only method has been applied to these datasets. This suggests that the audio stream is an important source of information that is often overlooked by prior work.
@@
Method Mod. Trim: Trim: Trim: Trim:
@@ 4 Modeling real data for deepfake detection
The previous section indicated that audio-visual datasets exhibit a silence bias which can easily be exploited. We want to develop a method that is robust to this bias (and possibly other uncovered ones) and still performs well. We have seen that models trained only on real data [17] are promising in being robust to the silence shortcut, but the performance was modest. To further improve them we propose to build on top of audio-focused self-supervised features. Self-supervised features have shown strong generalization for both visual [38, 13] and audio [40, 41] deepfake detection. We choose audio-focused self-supervised features because the audio models showed strong performance in the previous section. Note that we cannot rely on audio-only models because there are cases where manipulations appear only in the visual domain (the fake video, real audio case).
Figure 4: Overview of the AVH-Align method. A: We use the pretrained AV-HuBERT model to extract self-supervised features which we further align with a learnable network $\Phi$. Note that we use a single AV-HuBERT model, but make two forward passes to obtain audio-only and video-only features (instead of a single set of multimodal features). B: At training we maximize the alignment score $\Phi_{ii}$, between the audio features $\mathbf{a}_{i}$ at time step $i$ and the corresponding video features $\mathbf{v}_{i}$, while minimizing the alignment $\Phi_{ik}$ to the other features $\mathbf{v}_{k}$ in a neighboring window $\mathcal{N}(i)$.
@@ 4.1 Method
Self-supervised features. We use AV-HuBERT to represent both the audio and visual content of a video. AV-HuBERT is a Transformer network trained in a self-supervised way to predict iteratively refined centroids from masked inputs. The features extracted by AV-HuBERT encode audio information as proved by its strong performance on tasks such as lip reading or noisy audio-visual speech recognition. We extract the audio and visual representations independently: to extract audio features we mask the visual input, to extract visual features we mask the audio input. For a video, we obtain audio features $\mathbf{a}_{i}$ and visual features $\mathbf{v}_{i}$ features for each time step $i$. Both representations are 1024 dimensional and have a temporal resolution of 25 frames per second.
@@
The MLP has four layers, which progressively reduce the feature dimensionality, with layers mapping from the AV-HuBERT feature size of 1024 to 512, 256, 128, and finally to a single output. Each hidden layer includes Layer Normalization and ReLU activations.
@@
where $\mathcal{N}(i)$ represents the temporal neighborhood around the frame $i$. In our case $\mathcal{N}(i)$ contains the 30 neighboring frames around $i$. We define the final loss as the negative probabilities averaged across the entire video:
@@
This loss is similar to the contrastive loss InfoNCE [54], which was also used for deepfake detection [17, 39].
Inference. Once $\Phi$ is learned, we can estimate the fakeness score as the negative of $\Phi_{ii}$ for each audio–video frame pair in a video; aligned audio–video frame pairs should yield lower scores. Then we compute an overall alignment score for the entire video by pooling the per-frame scores using the log-sum-exp function (a smooth version of the max function).
Supervised variant. To understand the impact of the silence bias on the standard supervised learning paradigm, we design a supervised variant, AVH-Align/sup, that uses the same features and alignment network as AVH-Align, but a classification loss. In this setup we assume that apart from the real videos, we also have access to fake videos in the training set, with corresponding labels $y$. To obtain a per-video fakeness probability, we first pool the negated per-frame scores $\Phi_{ii}$ with the log-sum-exp function and then we apply the sigmoid function $\sigma$. Finally, we optimize the binary cross-entropy (BCE) loss:
$\mathcal{L}_{\text{sup}}=\text{BCE}\left(\sigma(\log\text{sum}\exp(-\Phi_{ii})),y\right).$ (4)
Metric: AUC Metric: AP
FakeAVCeleb AV-Deepfake1M FakeAVCeleb AV-Deepfake1M
Method Modality Train type Train data Trim: Trim: Trim: Trim: Trim: Trim: Trim: Trim:
@@
AVAD [17] AV unsup. LRS 84.5 84.7 $\cong$ 54.3 54.3 $\cong$ 99.5 99.5 $\cong$ 76.3 76.3 $\cong$
SpeechForensics [30] AV unsup. VoxCeleb2 98.8 98.8 $\cong$ 68.8 68.2 $\cong$ 100.0 100.0 $\cong$ 83.7 83.5 $\cong$
AVH-Align AV unsup. VoxCeleb2 94.6 94.6 $\cong$ 85.9 83.5 $\downarrow$ 99.8 99.8 $\cong$ 94.3 93.5 $\cong$
Table 2: Comparison of AVH-Align and AVH-Align/sup on FakeAVCeleb (FAVC) and AV-Deepfake1M (AV1M) original and trimmed datasets. AVH-Align is not impacted by the presence of leading silence in the original datasets, showing similar performance on the trimmed versions. In contrast, the AUC performance of AVH-Align/sup degrades by $16.9\%$ when tested on the trimmed variant of AV1M.
@@
Figure 5: Per frame fakeness probabilities for AVH-Align and AVH-Align/sup on AV-Deepfake1M. AVH-Align/sup always marks the first frame—corresponding usually to the leading silence—as fake, thus confirming that it uses the bias to distinguish between real and fake videos. AVH-Align is not affected by the presence of the leading silence. The fakeness probabilities for AVH-Align can be interpreted as misalignment probabilities, which is why they are higher during or after the manipulated region.